Some Trivia – the average volume of the driver side air bag in a regular car is ~56L while the passenger side air bag is 120L. Obviously these numbers vary a little base on the model and year of the car.

Question: If we were to use the reaction between Zn and HCl to fully inflate the driver side air bag (to 1 Atm) at 25oC, how many grams of Zn would be required.
(Assume you have sufficient HCl to react with this amount)
Answer #1

The equation for the reaction is:

Zn + 2 HCl \rightarrow ZnCl2 + H2

We use the ideal gas equation to determine the moles of H2.

moles (n) = PV / RT , where p is pressure, v is volume in liters, R is gas constant and t is temnperature in K. (298K)

n = 1 * 56 / (0.0821 * 298) =

2.289 moles

From the equation, moles of Zn needed = moles of H2 formed = 2.289 moles

Mass of Zn = moles * molar mass = 2.289 * 65.4 = 149.70 g